8) From the information given in the circuit below, (a) determine the current in the circuit, and (b) state which point, A or B, is at a higher potential. 27 Ω 30 V 8Ω 12Ω 10 V

answer is .38 A
Answer #1

Assume I current in the circuit is clock wise

KVL starting from a clockwise

30 - 27I - 10 - 12I - 8I - 5I = 0

I = 20/52 = 0.3846 Ampere

(b).

KVL from a to b

Va + 30 -I(27) = Vb

Va - Vb = (.38*27) - 30 = -19.6

Hence B is at higher potential then A.